I'm trying to find the indicated value of g(-12), g(-1), and g(1/4)the function given is g(x)=2x^2+5





Answer by Theo(13342)

that should not be hard to do.
I'll do one.
You can do the others.

your function is given as g(x) = 2x^2 + 5

g(-12) will therefore be equal to 2 * (-12)^2 + 5 which is equal to 2 * 144 + 5 which is equal to 288 + 5 which is equal to 293.

you can do the same with the other.

g(-1) = 2 * (-1)^2 + 5

g(1/4) = 2 * (1/4)^2 + 5

all you do is replace x with the number that replaced x as the argument.

g(x) = 2x^2 + 5 has x as the argument.

g(-12) replaces x as the argument with -12 to get:

g(-12) = 2 * (-12)^2 + 5